SPECTROSCOPY AND VIBRATIONAL PREDISSOCIATION OF Ar_{n}HF AT v_{HF} =3

\underline{CHENG-CHI CHUANG}, SUSY N. TSANG, JASON HANSON and WILLIAM KLEMPERER, Department of Chemistry, Harvard University, Cambridge, MA 02138; HUAN-CHNEG CHANG, Institute of Atomic and Molecular Sciences, Academica Sinica, P.O.Box 23-166, Taipei, Taiwan 10764, Republic of China.

Second overtone spectra of Ar_{n}HF (n=1,2 and 3) are observed by intracavity laser induced fluorescence. For ArHF(3000) the fluorescence radiation is dispersed and the rotational distribution of the HF (v=2) fragments are measured. The high intensity of fluorescence radiation is mainly due to the emission (v=3→1) of the undissociated ArHF, which indicates the lifetime of the (3000) state is significantly long, ≈ 2×10^{−5} sec. Additional new bend-stretch combination state of ArHF(3111) has been observed by both (3111)←(0100) hot band and (3111)←(0000) direct excitations. The preliminary spectroscopic constants of the (3111) are 11455.932 cm^{−1} and B=0.08490 cm^{−1} thus placing this state 116.898 cm^{−1} above the fundamental (3000). For Ar_{2}HF, four states correlating with j=0 and j=1 of HF(v=3) have been observed. The spectroscopic constants of these four bands are summarized as follows:

ν_{0}(band origin)  Δν  A  B  C
Σ_{0}  11323.783  -49.03 0.12023 0.05830 0.03894
Π_{in}  11387.730  63.95 0.12267 0.05705 0.03842
Π_{out}  11426.811  102.03 0.12026 0.05812 0.03863
Σ_{1}  11427.480  103.69 0.12026 0.05813 0.03764

The results provide direct probe of the three-body (non-additive) forces and are compared with the theoretical model of Ernesti and Hutson.^{a} For Ar_{3}HF the rovibrational lines of the v_{HF} =3 fundamental stretch state (ν_{0}=11310.451 cm^{−1}; 62 cm^{−1} red shift) are fitted by Voigt profile with Doppler and Lorenztian components of 100(10) and 50(10) MHz, respectively. The 50(10) MHz Lorenztian linewidths are attributed to the vibrational predissociation broadening (3 nsec in lifetime) for Ar_{3}HF at v_{HF} =3.

aA. Ernesti and J. M. Hutson, Phys. Rev. A51, 239 (1995).